OpenStudy (kittiwitti1):

A teapot is initially at rest on a horizontal tabletop, then one end of the table is lifted slightly. Does the normal force increase or decrease? [same question with] Force of static friction?

OpenStudy (kittiwitti1):

@AllTehMaffs

OpenStudy (kittiwitti1):

First part, yes; second part, idk

OpenStudy (kittiwitti1):

Actually I think it's yes for that too

OpenStudy (anonymous):

what do you mean by yes?

OpenStudy (kittiwitti1):

First part increases, static friction remains the same

OpenStudy (anonymous):

that's impossible.

OpenStudy (kittiwitti1):

Me or Y! Answers?

OpenStudy (anonymous):

\[ F_{static friction} = \mu_s N\]

OpenStudy (kittiwitti1):

It... looks like I'm wrong lols

OpenStudy (kittiwitti1):

Ok next question: Which is usually greater, the max force of static friction, or of kinetic friction?

OpenStudy (anonymous):

The normal force decreases, like you said (I didn't look at the yahoo thing). \[N = mgcos\theta\] If the pot is stationary the force of static friction though is equal to \[mg \sin\theta\] The coefficient of static friction would have to get much, much larger for those disparities. I dunno. tired.

OpenStudy (anonymous):

what's your guess on that last question?

OpenStudy (kittiwitti1):

Actually I can probably answer the conceptuals myself, but I have actual problems laters

OpenStudy (kittiwitti1):

So I'm good for the next 20 min or so...? No, make that 7 min

OpenStudy (anonymous):

k